# Towards a Generalized Theory of Observers

### v1

@@ 1 Introduction: Historical Perspectives on Observers
In quantum mechanics, the observer problem became explicit: measurement seemingly “collapses” the wavefunction, raising questions about consciousness, apparatus, and whether physical law alone dictates the outcome vonNeumann1955 ; Wheeler1989 . Meanwhile, in fields like cybernetics wiener1948 ; ashby1956 , biology maturana1980autopoiesis , and artificial intelligence newellSimon1976 , the observer is increasingly regarded as a feedback-driven agent or subsystem with the capacity to influence and be influenced by its environment.
@@ 10.2 Comparison with Existing Observer Theories
10.2.1 QBism: Formalizing Belief Updates with an Observer’s Boundary
@@
Quantum Bayesianism (QBism) regards the wavefunction as an agent’s subjective belief state about a quantum system, updating upon measurement. It emphasizes the personalist Bayesian perspective—probabilities (and quantum states) reflect an individual observer’s degrees of belief rather than objective system properties. QBism, however, treats the observer-agent as an axiomatic element, encoding interactions abstractly through broad axioms about rationality and coherence of beliefs. While QBism prescribes that observers update credences via the Born rule upon measurement, it does not provide a concrete operational mechanism describing how the belief states and the physical measurement process interrelate in practice. This leaves ambiguities: Where precisely is the interface between subjective experience and external quantum events? How exactly are belief updates physically realized?
Our minimal observer model explicitly addresses these ambiguities by providing observers with a concrete formal structure.
@@
In our minimal observer framework, an observer is explicitly defined by the tuple $O=(X,Y,Z,f,g,\mathcal{B})$. Here, $X$ is the observer’s internal state space (its memory or information states), and $\mathcal{B}$ is a well-defined boundary delineating the observer from the external environment. Measurement formally corresponds to an input $y\in Y$ (sensor reading) crossing the boundary $\mathcal{B}$ and subsequently updating the internal state via the transition function $f:X\times Y\to X$. Crucially, interpreting $f$ as a Bayesian update rule makes the QBist belief-update explicit and concrete. The observer’s internal state $x\in X$ formally represents their belief about the system, and when a measurement result $y$ is observed, $f(x,y)$ yields the updated belief state. The boundary $\mathcal{B}$ clearly defines the QBist observer-agent’s separation from the quantum system, resolving ambiguities inherent in QBist axioms by explicitly identifying the agent-system interface.
@@
Consider a QBist scenario involving a quantum coin-flip experiment. Initially, the observer has subjective credences (e.g., 50/50) for a biased quantum coin’s outcome. QBism merely states that the observer updates beliefs via the Born rule after observing an outcome. Our minimal observer model concretely depicts this update: the observer’s internal state $X$ initially encodes uncertainty. Upon measurement, the coin outcome crosses the boundary $\mathcal{B}$ into the observer’s sensory input space $Y$, leading $f$ to update the observer’s internal state $X$ with a definite belief state. Thus, the minimal observer model clarifies QBism’s subjective belief updates, making them tangible and precisely defined, thereby resolving ambiguities regarding the agent-system boundary and explicitly modeling the flow of information.
Furthermore, our model enables rigorous discussions of multiple observers with different priors, each updating their states consistently upon receiving measurements and communicating outcomes, thereby formalizing QBism’s inherently relational epistemic stance.
@@
Imagine two observers with distinct priors observing the same quantum coin toss. Each observer, represented formally by their own internal states $X$, independently updates their beliefs upon measurement. Only when they subsequently communicate (interactions crossing their respective boundaries $\mathcal{B}$) do their subjective beliefs align. Thus, our model operationally demonstrates QBism’s concept of observer-relative beliefs through explicit internal states and boundary interactions, enhancing interpretational clarity and potential experimental tests.

### v2

@@ 1 Introduction: Historical Perspectives on Observers
In quantum mechanics, the observer problem became explicit, ultimately raising questions about consciousness, apparatus, and whether physical law alone dictates the outcome of measurements vonNeumann1955 ; Wheeler1989 . As measurement seemingly “collapses” the wavefunction, intervention was recognized as a prerequisite of observation. In the field of cybernetics, a complementary development took place: observation was clearly recognized as a prerequisite for intervention, as observation (of outcomes) provides the input to adjust the actions of an agent in a feedback loop wiener1948 ; ashby1956 . A similar development can be observed in the fields of biology maturana1980autopoiesis , and artificial intelligence newellSimon1976 , where the observer is increasingly regarded as a feedback-driven agent or subsystem with the capacity to influence and be influenced by its environment.
@@ 10.2 Comparison with Existing Observer Theories
10.2.1 QBism: Formalizing Belief States with an Observer’s Boundary
@@
Quantum Bayesianism (QBism) treats the wavefunction as an observer’s subjective belief about a quantum system. Rather than describing objective system properties, quantum states reflect an agent’s personal degrees of belief. Importantly, QBism does not prescribe belief updates via the Born rule as probability updating (diachronic), but rather employs the Born rule to detect inconsistencies within an agent’s probability assignments at a given moment (synchronic coherence) Fuchs2023 . QBism similarly does not mandate Bayesian updating as the sole coherent strategy—belief updates depend fundamentally on each agent’s particular belief conditions Fuchs2023 ; DeBrota2018 .
While QBism clarifies measurement interactions conceptually, it deliberately avoids providing a third-person or physical dynamical mechanism explaining how belief states physically interact with measurement outcomes. This is intentional: QBism seeks to dissolve the measurement problem by reframing quantum mechanics as a decision-theoretic tool for agents, thus avoiding explanations via another physical process DeBrota2024 . Nevertheless, clarifying the agent–environment boundary remains conceptually beneficial, particularly in explicit modeling scenarios.
@@
Our minimal observer framework explicitly formalizes these concepts by defining an observer as a structured entity $O=(X,Y,Z,f,g,\mathcal{B})$, with internal states $X$ representing subjective beliefs, and boundary $\mathcal{B}$ explicitly distinguishing the observer from its environment. Measurement formally corresponds to sensory inputs $y\in Y$ crossing the boundary, updating the observer’s internal state via a general transition function $f:X\times Y\to X$. Crucially, $f$ need not represent solely Bayesian updating but can reflect any coherent belief-adjustment strategy consistent with the agent’s prior conditions. Thus, rather than resolving ambiguities in QBism, this formalism concretizes QBist principles in an operational manner, explicitly modeling the interface between subjective belief and external quantum events.
@@
Consider a quantum coin-flip scenario. Initially, an observer encodes subjective uncertainty regarding outcomes. Upon measurement, the observed outcome crosses the boundary $\mathcal{B}$ into sensory input space $Y$, prompting state-transition $f$ to yield a new internal belief state. This concrete depiction operationalizes QBism’s subjective belief perspective, explicitly modeling informational flow across agent-system boundaries. Such explicit modeling, while not required by QBism, offers clarity and enables exploration of scenarios involving multiple interacting observers.
@@
Imagine two observers with distinct priors observing the same quantum coin toss. Each observer, represented by internal state sets $X$, independently updates beliefs after measurements. Subsequent communication (interactions crossing boundaries $\mathcal{B}$) allows observers to exchange information; however, alignment of subjective beliefs is not guaranteed by mere communication alone, as Bayesian agents require substantial prior agreement (e.g., overlapping priors or common sample spaces) to reach consensus DeBrota2018 . This explicitly modeled observer framework thus allows rigorous and nuanced exploration of QBism’s relational epistemology, emphasizing the complexity and depth inherent in subjective belief alignment.
@@ 11 Conclusion and Outlook
Acknowledgments
We thank Stephen Wolfram for engaging discussions on observers. We also like to thank John DeBrota for useful suggestions and comments on QBism. This work was supported by the following grants: The Foundational Questions Institute and Fetzer-Franklin Fund, a donor advised fund of Silicon Valley Community Foundation [FQXi-RFP-1817]; The John Templeton Foundation [Grant ID 62106]; and the Australian Research Council [Grant DP210100919].